# APVR: Hour-Level Long Video Understanding with Adaptive Pivot Visual Information Retrieval

### v2

@@ 4.2 Comparison to the State-of-the-Art
Qualitative results
Fig. 3 presents representative qualitative results on hour-level video understanding, comparing our APVR framework with the baseline Qwen2.5-VL across four challenging samples. The results clearly demonstrate the unique advantages of APVR in both the semantic expansion and the subsequent retrieval process.
Firstly, the semantic information expansion module in APVR significantly enriches the original query, providing the downstream retrieval process with comprehensive object, attribute, and relational cues that are otherwise missing when using only the raw query. This richer semantic context enables the retrieval module to more accurately identify and prioritize frames that are truly relevant to the query intent.
Secondly, we observe that the pivot frames selected by APVR are highly correlated with both the query and the expanded semantic information, effectively filtering out irrelevant or redundant content. In long videos, there often exist temporally adjacent frames that are visually similar and all somewhat related to the query. If only a very limited number of frames are selected, it is easy to miss critical moments; conversely, retaining an excessive number of frames (e.g., 1024 frames) would cause out-of-memory (OOM) errors. APVR overcomes this dilemma by PFR and PTR and distills the most informative visual information within sufficient pivot frames. As a result, APVR empowers MLLM to efficiently comprehend long-form video content, providing accurate and detailed answers even in scenarios where baselines fail or generate hallucinations.
4.3 Ablation Studies
Ablation studies on PFR and PTR
To systematically validate the effectiveness of the proposed dual-granularity retrieval strategy, we conduct ablation studies on the two core components of APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTR), using the LongVideoBench benchmark. The results, shown in Fig. 4, consistently demonstrate that removing either PFR or PTR leads to a notable decline in accuracy across both baseline MLLMs. Without PFR, the system loses its ability to precisely localize query-relevant frames, resulting in performance degradation due to the inclusion of irrelevant or redundant content. Without PTR, the model is less able to compress visual tokens and focus on the most informative visual details, leading to processing maximum to only 256 frames and reduced answer quality. These findings strongly support the necessity and rationality of our dual-granularity pivot information retrieval design.
Ablation studies on other hyper-parameters
We further conduct a comprehensive analysis of key hyper-parameters that influence the performance of APVR. First, we examine the impact of the number of retrieved pivot frames $K$ and the weighting parameter $\lambda$ that balances the contributions of the two visual models in frame scoring. As summarized in Table 2, increasing the number of retained frames generally improves performance, especially on long videos, by reducing the risk of missing relevant segments. For the model weighting parameter $\lambda$, we observe that optimal results are achieved by striking a balance between CLIP-based semantic similarity and Grounding-DINO-based object detection, confirming the complementarity of these two perspectives.
Second, we analyze the effect of core retrieval hyper-parameters, specifically the number of retrieval iterations $p$ and the initial sampling stride $\nabla$. As shown in Table 3, the best performance is observed when $P=3$ and $\nabla=4$, indicating that a moderate number of adaptive iterations and a balanced initial sampling density allow the system to progressively refine frame selection without causing overfitting or excessive computational cost. Deviating from these settings leads to decreased accuracy, highlighting the importance of carefully tuning the retrieval process to maintain a balance between sampling and distribution update.
@@ 4.3 Ablation Studies
Table 2: Ablations of retained frame number and weight for scoring models.
$P$nabla Long Medium Short Average
3 2 56.4 62.1 77.3 62.9
3 4 58.3 65.8 77.3 64.9
3 8 55.1 64.6 77.3 63.1
2 4 56.0 66.0 77.3 63.9
4 4 58.0 65.8 77.3 64.7
5 4 58.5 64.5 77.3 64.5
Table 3: Ablations of iteration number and the initial sample stride.
5 Conclusion

### v3

@@ Introduction
Analysis on Information Retrieval
Fig. 3 presents representative qualitative results on hour-level video understanding, comparing our APVR framework with the baseline Qwen2.5-VL across four challenging samples. The results clearly demonstrate the advantages of APVR in both the semantic expansion and the subsequent retrieval process.
Firstly, we observe that the pivot frames selected by APVR are highly correlated with both the query and the expanded semantic information. APVR distills the most informative visual information within sufficient pivot frames. As a result, APVR empowers MLLM to efficiently comprehend long-form video content, providing accurate and detailed answers even in scenarios where baselines fail or generate hallucinations. Secondly, the semantic information expansion module in APVR significantly enriches the original query, providing the downstream retrieval process with comprehensive object, attribute, and relational cues that are otherwise missing when using only the raw query. As shown by ablation study results in Table 2, on LongVideoBench the performances decrease without expanded semantic information no matter the baseline MLLM.
@@
Table 3: Ablations of selected frames $K$ and weight $\lambda$.
Ablation Studies on PFR and PTR
To systematically validate the effectiveness of the proposed dual-granularity retrieval strategy, we conduct ablation studies for the two core components of APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTR), on LongVideoBench. The results, shown in Fig. 4, consistently demonstrate that removing either PFR or PTR leads to a notable decline in accuracy across all baseline MLLMs. Without PFR, the system loses its ability to precisely localize query-relevant frames, resulting in performance degradation due to the inclusion of irrelevant content (Fig. 1(a)). Without PTR, the model is less able to compress visual tokens and focus on the most informative visual details, leading to processing maximum to only 256 frames and reduced answer quality (Fig. 1(b) and (c)). These findings strongly support the necessity and rationality of our dual-granularity pivot information retrieval design.
Ablation Studies on Hyper-Parameters
We further conduct a comprehensive analysis of key hyper-parameters that influence the performance of APVR by ablation studies on LongVideoBench. First, we examine the impact of the number of retrieved pivot frames $K$ and the frame scoring weights $\lambda$. As shown in Table 3, increasing $K$ generally improves performance, especially on long videos, by reducing the risk of missing relevant segments. For $\lambda$, we observe that optimal results are achieved by striking a balance between spatial and semantic scoring, confirming the complementarity of these two perspectives.
Second, we analyze the effect of the number of frame retrieval iterations $p$ and the initial sampling stride $\nabla$. As shown in Table 4, the best performance is observed when $P=3$ and $\nabla=4$, indicating that a moderate number of adaptive iterations and a balanced initial sampling density allow the system to progressively refine frame selection without causing overfitting or excessive computational cost.
$P$ $\nabla$ Long(10-60m) Medium(1-10m) Average
3 2 56.4 62.1 62.9
3 4 58.3 65.8 64.9
3 8 55.1 64.6 63.1
2 4 56.0 66.0 63.9
4 4 58.0 65.8 64.7
5 4 58.5 64.5 64.5
Table 4: Ablations of iteration number $P$ and stride $\nabla$.
Conclusion
@@
Appendix
Prompt Template for Query Expansion
The prompt in Fig. 5 was used to interact with the language model for semantic information expansion.
Query Expansion Prompt Example Analyze the following video understanding question: Question: <Question>; Options: <Options> Step 1: Key Object Identification • Extract 3-5 core objects detectable by computer vision • Use Grounding dino compatible noun phrases (e.g., “person”, “mic”) • Format: Key Objects: obj1, obj2, obj3 Step 2: Contextual Cues • List 2-4 scene elements that help locate key objects based on options provided • Use detectable items (avoid abstract concepts) • Format: Cue Objects: cue1, cue2, cue3 Step 3: Relationship Triplets • Relationship types: • Spatial: Objects must appear in the same frame • Attribute: Color/size/material descriptions (e.g., “red clothes”, “large”) • Time: Appear in different frames within a few seconds • Causal: There is a temporal order between the objects • Format of Relations: (object, relation type, object), relation type should be exactly one of spatial/attribute/time/causal Step 4: Description Augmentation • List 1-3 descriptions based on knowledge graph augmentation for each object • Entity descriptions (e.g., “mic is a device for amplifying sound”) • Hypernym concepts (e.g., “dog is a kind of animal”) • Format: Description: (object: des1; des2) Step 5: Semantics Augmentation • List 2-5 Semantics information of query and options based on knowledge graph (e.g., “leash often appears with dog”) • Format: Semantics: semantic1; semantic2 Output Rules 1. One line each for Key Objects/Cue Objects/Relation/Des/Sem starting with exact prefixes 2. Separate items with comma except for triplets where items are separated by semicolon 3. Never use markdown or natural language explanations 4. If you cannot identify any key objects or cue objects from the video provided, please just identify the possible key or cue objects from the question and options provided Below is an example of the procedure: Question: For “When does the person in red clothes appear with the dog?” Response: Key Objects: person, dog, red clothes Cue Objects: grassy area, leash, fence Rel: (person; attribute; red clothes), (person; spatial; dog) Des: (red clothes: description1), (dog: description2) Sem: semantic1; semantic2 Format your response EXACTLY like this in Five lines: Key Objects: object1, object2 Cue Objects: object3, object4 Rel: (object1; relation type1; object2), (object3; relation type2; object4), … Des: (object1: description1; description2), (object2: description1; description2), … Sem: semantic1; semantic2, …
Figure 5: The Prompt Template for Quey Expansion.
The Details of Spatio-semantic Confidence Scoring
Semantic Scoring
For each of the sampled frames $\mathcal{F}_{samp}=\{f_{t}\}_{t=1}^{N_{s}}$, we encode it using the CLIP image encoder $\phi_{I}(\cdot)$:
[…]
The query as well as the expanded description and semantics are aggregated and encoded using the CLIP text encoder $\phi_{T}(\cdot)$:
[…]
The CLIP similarity score for frame $f_{t}$ is then calculated as:
$s_{t}^{CLIP}=softmax(\tau\cdot v_{t}\cdot t_{agg}),$ (19)
where $\tau$ is a temperature parameter set to $100$ to sharpen the score distribution.
Spatial Scoring
To enhance the retrieval precision with key/cue objects along with their relations, we first employ Grounding-DINO to identify extracted objects:
$\mathcal{B}_{t},\mathcal{L}_{t}=\Psi_{GD}(f_{t},O_{k}\cup O_{c}),$ (20)
where $\Psi(\cdot)$ is the Grounding-DINO model that returns bounding boxes $\mathcal{B}_{t}$ and corresponding confidence logits $\mathcal{L}_{t}$ for frame $f_{t}$. The score of detection is then computed as:
$s_{t}^{o}=softmax(max(\mathcal{L}_{t})).$ (21)
To further model the logic relations among objects in the same or different frames to enhance the detection results, we update the score by adding corresponding frame score if there are satisfied relation triplets in this frame as follows:
$s_{t}^{GD}=s_{t}^{o}+\sum_{r}(\omega_{r}\cdot s_{r}),$ (22)
where $r$ is exactly one of spatial/time/attribute/causal and $\omega_{r}$ is the corresponding weight to control the contribution of each type of relation.